# Every Wrinkle Carries A Memory: An Integro-differential Bootstrap for Features in Cosmological Correlators

### v1

@@ 5.1 IR resonances
giving rise to a non-trivial constraint equation for $\alpha$, [^31] Notice that the dependence on $\eta_{0}$ miraculously cancel out in the characteristic equation, showing that the resonance is not sensitive to the phase of mass oscillations.
@@ 5.2 UV resonances
$\Delta\beta_{-}$ $\approx-\dfrac{i\alpha_{0}^{*}\,g^{2}m_{0}^{2}}{4}\int^{s\eta_{\rm end}}_{-\infty}\dfrac{dx}{x^{2}}e^{2ix}\,(-x)^{+i\omega}\approx-2\sqrt{\pi}g^{2}m_{0}^{2}\alpha_{0}^{*}\,e^{-i(\omega+\pi/4)}\left(\frac{\omega}{2}\right)^{-\frac{3}{2}+i\omega}\,$, (5.46)
@@ 6.2 Results
Figure 10: Convergence tests of the numerical bootstrap solution with respect to the mesh points NN (left panel) and the cutoff $L=\log(k_{12}/s)_{\text{max}}$ (right panel). The benchmark kinematics is chosen to be $r_{*}=\log(k_{12}/s)_{*}=8$ and the blue and yellow points correspond to $g=0.2$ and $g=0.3$, respectively. The dashed lines represent the prediction of the perturbative analytical solution. The other parameters are chosen as $\mu=1$, $\omega=1/3$, $x_{0}=-s\eta_{0}=2$. The boundary conditions are implemented at $r_{\rm I}=1$ and at $r_{\rm II}=1+2L/N$. The regularisation gap is chosen to be $\delta=10^{-1}$. In the left panel, the kinematic cutoff is fixed to be $L=20$ and in the right panel, the mesh points are chosen such that the step size $L/N=10^{-2}$ is fixed. We conclude that the numerical solution converges at large $N$ and $L$ towards the perturbative prediction, but with a small $\mathcal{O}(g^{4})$ mismatch. The weak dependence on $L\gtrsim r_{*}$ also shows the quasi-locality of the bootstrap equation.
@@
To test the universality of the scaling exponent and its independence of the boundary conditions, we explore the collective behaviour of solutions by randomly sampling boundary conditions in the vicinity of our perturbative Bunch-Davies solution. This can be achieved via deforming the boundary conditions $\mathcal{C}_{\rm b.c.}\mathbf{F}=Z\bm{\mathcal{S}}_{\rm b.c.}$ by a random complex matrix $Z$. We implement this by modifying $F(r_{\rm I})=Z_{\rm I}F^{(g^{2})}(r_{\rm I})$ and $F(r_{\rm II})=Z_{\rm II}F^{(g^{2})}(r_{\rm II})$ with $Z_{\rm I},Z_{\rm II}$ two random complex numbers generated from a log-normal distribution $\exp(-\frac{1}{2}\log^{2}|Z|)$. In Figure 13, we plot the scaling behaviour of the cosmological collider signal and confirm that it is indeed independent of the boundary conditions and that it agrees with the prediction from the eigenfrequency analysis in Section 5.1,
@@ 7  Summary and outlooks
  • Dispersive bootstrap: As we have seen in Section 4.2, the factorisation property of the exchange diagram imposes highly non-trivial constraints on its analytic structure. Along the lines of [75, 33], this suggests an alternative bootstrap approach, in which the exchange graph—without invoking the inhomogeneous IDE—is computed directly from the knowledge of the three-point function $f$, using an appropriate dispersive representation in the $s^{2}$ complex plane. By following this method it might be possible to obtain different analytic expressions for the exchange diagram with better convergence properties near the equilateral limit $U=1$, relative to our series expansion formulas.
@@ Acknowledgments
We would like to thank Ali Akbar Abolhasani, Xingang Chen, Siyao Li, Zhehan Qin, Sébastien Renaux-Petel, Enrico Pajer, Guilherme L. Pimentel, Arthur Poisson, David Stefanyszyn, Denis Werth, Dong-Gang Wang, Zhong-Zhi Xianyu and Bowei Zhang for stimulating discussions. We thank Xingang Chen, Guilherme L. Pimentel, Lucas Pinol and Sébastien Renaux-Petel for comments on the manuscript. SJ would like to thank the organisers of the workshop “Constraining Effective Field Theories without Lorentz” (July 2025 at IFPU, Trieste) for their hospitality, when part of this work was under development, and the participants for lively discussions. SJ is supported by Simons Investigator Award 690508. X.T. is supported by STFC consolidated grant ST/X000664/1. Y.Z. is supported by the IBS under the project code, IBS-R018-D3.
@@ Appendix A An infinite set of coupled bootstrap equations
Up to ${\cal O}(g^{4})$ corrections to $F_{++}^{(0,0)}$, the last two equations can be solved by ignoring the would-be sources on the right-hand side. Plugging the solutions for $F^{(\pm 1,0)}_{++}$ back into the first line, and after decomposing $F^{(0,0)}_{++}$ as (4.35), it is straightforward to show that the right-hand sides of (A.8) and (4.37) will be equal, consistent with the fact that the two sets of equations describe identical correlators.

### v2

@@ 5.1 IR resonances
giving rise to a non-trivial constraint equation for $\alpha$,
@@
Notice that the dependence on $\eta_{0}$ miraculously cancel out in the characteristic equation, showing that the resonance is not sensitive to the phase of mass oscillations.
@@ 5.2 UV resonances
$\Delta\beta_{-}$ $\approx-\dfrac{i\alpha_{0}^{*}\,g^{2}m_{0}^{2}}{4}\int^{s\eta_{\rm end}}_{-\infty}\dfrac{dx}{x^{2}}e^{2ix}\,(-x)^{+i\omega}\approx-i\sqrt{\frac{\pi}{2}}g^{2}m_{0}^{2}\alpha_{0}^{*}\,e^{-i(\omega+\pi/4)}{\omega}^{-\frac{3}{2}+i\omega}\,2^{-i\omega}\,$, (5.46)
@@ 6.2 Results
Figure 10: Convergence tests of the numerical bootstrap solution with respect to the mesh points NN (left panel) and the cutoff $L=\log(k_{12}/s)_{\text{max}}$ (right panel). The benchmark kinematics is chosen to be $r_{*}=\log(k_{12}/s)_{*}=8$ and the blue and yellow points correspond to $g=0.2$ and $g=0.3$, respectively. The dashed lines represent the prediction of the perturbative analytical solution. The other parameters are chosen as $\mu=1$, $\omega=1/3$, $x_{0}=-s\eta_{0}=1$. The boundary conditions are implemented at $r_{\rm I}=1$ and at $r_{\rm II}=1+2L/N$. The regularisation gap is chosen to be $\delta=10^{-1}$. In the left panel, the kinematic cutoff is fixed to be $L=20$ and in the right panel, the mesh points are chosen such that the step size $L/N=10^{-2}$ is fixed. We conclude that the numerical solution converges at large $N$ and $L$ towards the perturbative prediction, but with a small $\mathcal{O}(g^{4})$ mismatch. The weak dependence on $L\gtrsim r_{*}$ also shows the quasi-locality of the bootstrap equation.
@@
To test the universality of the scaling exponent and its independence on the boundary conditions, we explore the collective behaviour of solutions by randomly sampling boundary conditions in the vicinity of our perturbative Bunch-Davies solution. This can be achieved via deforming the boundary conditions $\mathcal{C}_{\rm b.c.}\mathbf{F}=Z\bm{\mathcal{S}}_{\rm b.c.}$ by a random complex matrix $Z$. We implement this by modifying $F(r_{\rm I})=Z_{\rm I}F^{(g^{2})}(r_{\rm I})$ and $F(r_{\rm II})=Z_{\rm II}F^{(g^{2})}(r_{\rm II})$ with $Z_{\rm I},Z_{\rm II}$ two random complex numbers generated from a log-normal distribution $\exp(-\frac{1}{2}\log^{2}|Z|)$. In Figure 13, we plot the scaling behaviour of the cosmological collider signal and confirm that it is indeed independent of the boundary conditions and that it agrees with the prediction from the eigenfrequency analysis in Section 5.1,
@@ 7  Summary and outlooks
  • Dispersive bootstrap: As we have seen in Section 4.2, the factorisation property of the exchange diagram imposes highly non-trivial constraints on its analytic structure. Along the lines of [76, 33], this suggests an alternative bootstrap approach, in which the exchange graph—without invoking the inhomogeneous IDE—is computed directly from the knowledge of the three-point function $f$, using an appropriate dispersive representation in the $s^{2}$ complex plane. By following this method, it might be possible to obtain alternative analytic expressions for the exchange diagram with better convergence properties near the equilateral limit $U=1$ than our series expansion formulas.
@@ Acknowledgments
We would like to thank Ali Akbar Abolhasani, Sebastián Céspedes, Xingang Chen, Siyao Li, Zhehan Qin, Sébastien Renaux-Petel, Guilherme L. Pimentel, Arthur Poisson, David Stefanyszyn, Denis Werth, Dong-Gang Wang, Zhong-Zhi Xianyu and Bowei Zhang for stimulating discussions. We thank Xingang Chen, Guilherme L. Pimentel, Lucas Pinol and Sébastien Renaux-Petel for comments on the manuscript. SJ would like to thank the organisers of the workshop “Constraining Effective Field Theories without Lorentz” (July 2025 at IFPU, Trieste) for their hospitality, when part of this work was under development, and the participants for lively discussions. S.J. is supported by the STFC Consolidated Grant ST/X000575/1 and a Simons Investigator Award 690508. X.T. is supported by STFC consolidated grants ST/T000694/1 and ST/X000664/1. Y.Z. is supported by the IBS under the project code, IBS-R018-D3.
@@ Appendix A An infinite set of coupled bootstrap equations
Up to ${\cal O}(g^{4})$ corrections to $F_{++}^{(0,0)}$, the last two equations can be solved by ignoring the would-be sources on the right-hand side. Plugging the solutions for $F^{(\pm 1,0)}_{++}$ back into the first line, and after decomposing $F^{(0,0)}_{++}$ as (4.35), it is straightforward to show that the right-hand sides of (A.8) and (4.37) will match, consistent with the fact that the two sets of equations describe identical correlators.
@@ B.2 Three-point exchange diagram
Appendix C Light fields and the anomalous scaling
We have mostly been concerned with heavy field exchanges in the main body of this work. However, the analytical and numerical techniques readily extend to the case of light fields. First, we note that the boundary bootstrap equation retains its form (1.10), albeit with a light mass $m_{0}$ lying in the complementary series with $\nu=(9/4-m_{0}^{2}/H^{2})^{1/2}>0$. To avoid IR divergences in the $\dot{\pi}_{c}\sigma$ and $\dot{\pi}_{c}^{2}\sigma$ vertices. In the absence of parametric resonances this implies
$\int{\rm d}\eta\,a^{3}(\eta)\pi_{c}^{\prime}(\eta)\sigma(\eta)\sim\int{\rm d}\eta\,(-\eta)^{-1/2-\nu}\quad\Rightarrow\quad 0<\nu<1/2~,$ (C.1)
meaning that the $\sigma$ field must be heavier than the conformally coupled scalar. [^21] In the scale-invariant case $g=0$, the IR divergences in the Schwinger-Keldysh diagrams $F_{\sf ab}$ are known to cancel out in correlator observables [45]. In the scale-breaking case $g\neq 0$, it remains unclear if the cancellation persists, we therefore choose to assume a sufficiently large mass to ensure IR convergence. In this regime, the perturbative solution in (4.64) and (4.85) directly translates to the light-mass case after replacing $\mu\to-i\nu$, resulting in an analytical solution accurate at the order of $g^{2}$. One might therefore expect the translation to the light-mass case to be trivial at each order. Yet surprisingly, this turns out not to be the case. In fact, we discover an entirely new effect emerging at the order of $g^{4}$, which will first manifest itself in the numerics below.
Numerical bootstrap
The numerical bootstrap can be performed in the same fashion as in the heavy-mass case. We confirm the numerical convergence in Figure 14. From the plots, we also observe that the non-locality in the light-field-exchange case appears to be much stronger than in the heavy one (c.f. Figure 10), as the numerics only converges for $L-r_{*}=\mathcal{O}(10)$. This matches the intuition that light fields propagate a longer distance than heavy fields and are therefore more non-local.
Figure 14: Convergence tests of the numerical bootstrap solution for the light-field exchange with respect to the mesh points NN (left panel) and the cutoff $L=\log(k_{12}/s)_{\text{max}}$ (right panel). The benchmark kinematics is chosen to be $r_{*}=\log(k_{12}/s)_{*}=8$ and the blue and yellow points correspond to $g=0.2$ and $g=0.3$, respectively. The dashed lines represent the prediction of the perturbative analytical solution. The other parameters are chosen as $\nu=0.2$, $\omega=1/3$, $x_{0}=-s\eta_{0}=1$. The boundary conditions are implemented at $r_{\rm I}=1$ and at $r_{\rm II}=1+2L/N$. The regularisation gap is chosen to be $\delta=10^{-1}$. In the left panel, the kinematic cutoff is fixed to be $L=20$ and in the right panel, the mesh points are chosen such that the step size $L/N=10^{-2}$ is fixed. Thus the numerical solution again converges at large $N$ and $L$ towards the perturbative prediction. Comparison to the heavy field case in Figure 10, we observe a much stronger non-locality of the bootstrap equation.
The waveforms normalised according to the naive scaling exponent $1/2+\nu$ are shown in Figure 15. Apparently, the numerical solutions start to deviate from the naive scaling at large couplings $g$ and in the squeezed limit $k_{12}/s\gg 1$, signalling the breakdown of the $\mathcal{O}(g^{2})$ perturbative solution. Carefully examining the plots indeed points to the emergence of a new expansion parameter $g^{4}\log(k_{12}/s)$ which is absent in our analytical solution at order $\mathcal{O}(g^{2})$ due to its high order in $g$. A resummation of this new expansion parameter would then lead to an anomalous correction to the scaling exponent in the squeezed limit:
$F(k_{12},s)\sim\left(\frac{k_{12}}{s}\right)^{-1/2+\nu}\left(1+\mathcal{O}(g^{4})\times\log\frac{k_{12}}{s}+\cdots\right)\to\left(\frac{k_{12}}{s}\right)^{-1/2+\nu+\mathcal{O}(g^{4})}~.$ (C.2)
Figure 15: The numerical solution (solid curves) as a function of the momentum ratio $r=\log(k_{12}/s)$ normalised by the naive scaling exponent. The dotted curves indicates the anomalous scaling predicted from the boundary (C.4). The left and right panels correspond to frequency choices $\omega=2/3$ and $\omega=4/3$, respectively. The grey, blue and yellow curves correspond to different sizes of the coupling $g$ as specified in the right panel. Other model parameters are chosen as $\nu=0.2$, $x_{0}=-s\eta_{0}=2$. The number of mesh point is $N=10^{4}$ with a kinematic cutoff $L=40$. The boundary conditions are implemented at $r_{\rm I}=1$ and at $r_{\rm II}=1+2L/N$ with a regularisation gap $\delta=10^{-1}$.
Anomalous scaling from the boundary
Interestingly, the emergence of a new scaling can be very easily predicted from the boundary. In fact, the anomalous scaling exponent can be analytically computed via the same eigenfrequency analysis as in the IR resonances. We make the same ansatz as in (5.7),
$F(k_{12},s)=\sum_{n=-\infty}^{\infty}\xi_{n}\left(\frac{k_{12}}{s}\right)^{\alpha+in\omega}~,$ (C.3)
and truncate the resultant recurrence relation at $\xi_{\pm 2}=\xi_{\pm 3}=\cdots=0$. The constraint equation is identical to (5.17) but with $\mu\to-i\nu$. Now, instead of expanding around $\omega=2\mu$, we directly solve the constraint near $\alpha=-1/2+\nu$, yielding {keyeqn}
$\alpha=-1/2+\nu-\frac{g^{4}\left(9/4-\nu^{2}\right)^{2}}{4\nu\left(4\nu^{2}+\omega^{2}\right)}+\mathcal{O}(g^{8})~,$ (C.4)
which produces exactly the expected scaling exponent (see the dotted lines in Figure 15). This eigenfrequency analysis therefore non-perturbatively resums large logarithms of the momentum ratio from a pure boundary perspective.
Effective mass and an analogy to Kapitza’s pendulum
$\textsl{b}\cos(\omega t)\ellmmg$$\theta$
Figure 16: Kapitza’s pendulum of mass $m$, length $\ell$, frequency $\omega$, amplitude b and gravity g.
The anomalous scaling can be understood through the analogy to a classical mechanics system known as the Kapitza pendulum [131]. The Kapitza pendulum consists of an inverted pendulum with a vertically oscillating pivot point (see Figure 16 for an illustration). The nature of the inverted position implies an instability for the pendulum, yet the fast-oscillating pivot point can distort its motion and effectively stabilise it. Up to total derivatives, its Lagrangian
$L=\frac{1}{2}m\ell^{2}\dot{\theta}^{2}-m\ell\left[\textsl{g}-\textsl{b}\,\omega^{2}\cos(\omega t)\right]\cos\theta$ (C.5)
leads to an equation of motion of the form
$\left[\frac{\partial^{2}}{\partial t^{2}}-\frac{\textsl{g}}{\ell}\left(1-\frac{\textsl{b}\,\omega^{2}}{\textsl{g}}\cos(\omega t)\right)\right]\hat{\theta}+\mathcal{O}(\hat{\theta}^{3})=0~,$ (C.6)
where we have performed a linear expansion around $\theta=0$ and $\hat{\theta}\equiv\sqrt{m}\ell\theta$ is the canonical variable. In the double-scaling limit $\textsl{b}\to 0$, $\omega\to\infty$ with $\textsl{b}\,\omega$ fixed, perturbative analysis using the separation of scales shows that the pendulum oscillates with an effective potential [132]
$\frac{\partial^{2}\hat{\theta}}{\partial t^{2}}=-\frac{\partial V_{\rm eff}(\hat{\theta})}{\partial\hat{\theta}}~,\quad V_{\rm eff}(\hat{\theta})=\frac{1}{2}\left(-\frac{\textsl{g}}{\ell}+\frac{\textsl{b}^{2}\omega^{2}}{2\ell^{2}}\right)\hat{\theta}^{2}+\mathcal{O}(\hat{\theta}^{4})~.$ (C.7)
The second term in the last bracket provides a positive correction to the effective frequency of the pendulum which tends to stabilise the system. Comparing (C.6) to the equation of motion of the $\sigma$ field with light masses,
$\left[\frac{\partial^{2}}{\partial t^{2}}-\nu^{2}\left(1-\tilde{g}^{2}\cos(\omega t)\right)\right](a^{3/2}\sigma(s,t))\approx 0~,\quad\tilde{g}^{2}\equiv g^{2}\left(\frac{9}{4\nu^{2}}-1\right)~,$ (C.8)
we make the identification
$\nu^{2}=\frac{\textsl{g}}{\ell}\quad,\quad\tilde{g}^{2}=\frac{\textsl{b}\,\omega^{2}}{\textsl{g}}~.$ (C.9)
Thus the corrected effective mass reads
$\nu_{\rm eff}^{2}=\nu^{2}-\frac{\tilde{g}^{4}\nu^{4}}{2\omega^{2}}~,$ (C.10)
leading to the effective scaling exponent
$(\alpha)_{\rm Kapitza}=-1/2+\nu_{\rm eff}=-1/2+\nu-\frac{g^{4}\left(9/4-\nu^{2}\right)^{2}}{4\nu\omega^{2}}~.$ (C.11)
This matches nicely with our result from the boundary bootstrap (C.4) in the double-scaling limit with $\omega\gg\nu$. Taking the analogy, we therefore conclude that time-dependent oscillations provide an (necessarily positive) effective correction to the mass of the light field, resulting in the anomalous scaling in the squeezed limit of boundary correlators. It is now transparent why this effect only appears at order $\mathcal{O}(g^{4})$ since one needs at least two insertions of cosine oscillations to generate a non-zero effective mean i.e. $\overline{\cos^{2}(\omega t)}=1/2$. This also suggests that such corrections exist for heavy fields as well. A complete understanding of this phenomenon would require an analytical investigation at order $\mathcal{O}(g^{4})$. We leave such a more thorough analysis for future works.